solve for x
2=4+2x

Answer by jim_thompson5910(35256)

2 = 4 + 2x

2 - 4 = 4 + 2x - 4

-2 = 2x

-2/2 = 2x/2

-1 = x

x = -1


So the solution is x = - 1